2. Find the total surfa,ce area of a cone, if its slant height is
21 m and diameter of its ba,se is 24 m.

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
lbtal surface area = rr:rl + rr:r2 = nr( l + r) = 722" x 12 x (21 + 12) cm2 = 1244.57 m2•